OpenStudy (anonymous):

urgent please help! write the standard form of a complex number giving two examples

OpenStudy (anonymous):

any complex number?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i vote for \(2+3i\) but you can make up your own

OpenStudy (anonymous):

is that the complex or standard version?

OpenStudy (anonymous):

i sense some confusion here

OpenStudy (anonymous):

there is no such thing as "complex" or "standard" version there is "standard form" for a complex number, which is \(a+bi\)

OpenStudy (anonymous):

so for example \(2+3i\) is in standard form, but \[\frac{2+3i}{5}\] is not

OpenStudy (anonymous):

\(\frac{1+\sqrt{3}i}{2}\) is not standard form, but \(\frac{1}{2}+\frac{\sqrt{3}}{2}i\) is

OpenStudy (unklerhaukus):

A complex number is a number that has both a real part and an imaginary part. If \(z\) is a complex number ; it can be written in this form \[z=a+ib\] where \(a\) and \(b\) are both real numbers \(a={\frak{R}}( z) \) is the real part of \(z\) and \(b={\frak{I}}( z) \) is the imaginary part of \(z\) When the number has no imaginary component \(b=0\) and we call this number a real number \(z\in \mathbb R\). When the number has no real component \(a=0\) and we call this number a purely imaginary number \(z\in \mathbb I\). When neither \(a\) nor \(b\) are zero the number has both a real part and an imaginary part. Only these numbers are complex \(z\in \mathbb C\) So to construct a complex number just choose two non zero real numbers, multiply one by the imaginary constant \(i\), and add them.